OpenStudy (anonymous):

Which of the following best describes the technique used to graph the equation, y = -2x + 8, using slope and y-intercept? Mark a point at 8 on the y-axis, move upward 2 and right 1, mark the point and draw the line. Mark a point at 8 on the y-axis, move downward 2 and right 1, mark the point and draw the line. Mark a point at -2 on the y-axis, move downward 8 and right 1, mark the point and draw the line. Mark a point at -2 on the y-axis, move upward 8 and right 1, mark the point and draw the line.

OpenStudy (anonymous):

True or False? Y varies directly with x and y = -9 when x = 3. When y = 27, the value of x is 9.

OpenStudy (michele_laino):

when x=3 then y is: y=-2*3+8=+2

OpenStudy (michele_laino):

I used your equation y=-2x+8

OpenStudy (dtan5457):

y = -2x + 8 8 is your intercept, and it has a negative slope of 2. Which above describes that?

OpenStudy (anonymous):

WOW

OpenStudy (anonymous):

so its true for the second question?

OpenStudy (dtan5457):

Using the original function: y = -2x + 8 "Y varies directly with x and y = -9 when x = 3. When y = 27, the value of x is 9." Plug them in and you tell me. I'll plug the first pair in. -9=-2(3)+8 6+8=14, that's false. Try the 2nd one.

OpenStudy (anonymous):

-9=-2(3)+8 is false

OpenStudy (dtan5457):

Yes ^